. Find the number of years it would take for R748,00

to increase by 72% if it is invested at an interest rate of 7,9% per annum, compounded semi-annually


1
Expert's answer
2021-06-30T10:10:40-0400

Amount, "A=748\times 172/100=R1286.56"

"A=P(1+r/100)^{2n}, r=7.9/2=3.95\%"

"1286.56=748(1+3.95\/100)^{2n}"

"1.72=(1.0395)^{2n}"

"(log1.72)\/(log1.0395)=2n"

"2n=13.999"

"n=7yrs"